Cost of Units Completed and in Process

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 2,000 units, 30% completed 5,540 To Finished Goods, 46,000 units ?
Direct materials, 47,000 units @ $1.6 75,200
Direct labor 142,300
Factory overhead 55,310
Bal. ? units, 55% completed ?

a. Based on the above data, determine the different costs listed below.

If required, round your interim calculations to two decimal places.

1. Cost of beginning work in process inventory completed this period. $
2. Cost of units transferred to finished goods during the period. $
3. Cost of ending work in process inventory. $
4. Cost per unit of the completed beginning work in process inventory, rounded to the nearest cent. $

b. Did the production costs change from the preceding period?

c. Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?
Answer #1
Ending Inventory
Beginning Inventory 2000
Add: Units placed in production 47000
Less: Finished Goods -46000
Ending Inventory 3000
Equivalent Units
Equivalent Units
Whole Units Direct Material Conversion
Beginning Inventory (30% completed)=(2000*(1-30%)) 2000 1400
Started and completed(46000-2000) 44000 44000 44000
Transferred to finished goods 46000 44000 45400
Ending Inventory(45% completed)=(3000*55%) 3000 3000 1650
Total Units 49000 47000 47050
Cost
Direct Material Conversion
Total cost for the Assembly Department=(A) $         75,200.00 $     1,97,610.00
Total Equivalent units=(B) 47000 47050
Cost per Equivalent Units=(A)/(B) $                   1.60 $                    4.20
Conversion cost=($142300+$55310)
Cost of units started and completed=(44000*($1.60+$4.20)) $     2,55,200.00
a)
1) Beginning Work in Process balance(Given) $           5,540.00
Cost of conversion cost during the period=(1400*$4.20) $           5,880.00
Cost of beginning inventory $         11,420.00
2) Cost of beginnning Work in Process $         11,420.00
Cost of units started and completed during period $     2,55,200.00
Cost of units transferred to finished goods $     2,66,620.00
3) Cost of Ending Work in Process
Direct Material (Equivalent units* Cost per equivalent units)=(3000*$1.60) $           4,800.00
Conversion (Equivalent units*Cost per equivalent units)=(1650*$4.20) $           6,930.00
Cost of Ending Work in Process $         11,730.00
4) Cost per unit of beginning work in process inventory=($11420/2000) $                   5.71
b) Yes the production cost changed during the current period.
Earlier Period Current Period
Cost of production 5.71 $                    5.80
c) Beginning Work in Process(given) $           5,540.00
Less: Direct Material incurred in the previous period(2000*$1.60) $         -3,200.00
Conversion cost incurred in the previous period $           2,340.00
Conversion cost per unit in the current period $                   4.20
Conversion cost per unit in previous period=($2340/(2000*30%)) $                   3.90
Increase conversion cost per equivalent units $                   0.30